OpenStudy (anonymous):

Over the last decade, the population of Merrifield has increased by 110%. If the population was 34,450 a decade ago, which number is closest to the population today?..

OpenStudy (anonymous):

I TIMED 34450 TO 110% THEN ADD IT TO 34450 I KEEP GETTING A BIG NUMBER

OpenStudy (anonymous):

ok its like this

OpenStudy (anonymous):

110% ---- =---- 100 34450

OpenStudy (anonymous):

37895? is the population now?

OpenStudy (anonymous):

so 110 * 34450 =3,789,500 then u divide that number by 100 which is 3,789,500/100=37,895 Yes that is your answer

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

34450+ 34450= 68900 100% of 34450 is itself then u got 10% left over . 34450/10 is 3445 then u add 3445 to 63900 3445+63900=67345 so i think 67345 could be also an answer if the other one is wrong. i think its 67345. thats your answer lol

OpenStudy (anonymous):

oh my mbad the answer is 68900+3445 which is 72345

OpenStudy (anonymous):

that is surely your answer

OpenStudy (anonymous):

i looked at 68900 wrong lol but 72345 is still the right answer